# A geometric approach to exponentially small splitting: The generic zero-Hopf bifurcation of co-dimension two

### 12 Mar 2026 v1

@@ 5. A dynamical systems approach to the difference
The system (5.14) can be viewed as a Fenichel normal form, see [25, Section 3]. The following result shows that the diagonalization (5.13) can be “globalized” beyond a compact subset $v_{2}\in I$ (by working in the directional charts $\breve{v}=\pm 1$ associated with the blowup (5.4)).
@@
Now, notice that with respect to the backward flow, the center manifolds in the $\breve{v}=\pm 1$-charts and the slow manifold in the $\breve{\epsilon}=1$-chart all become attracting and moreover $v$ (in particular $v_{2}$) becomes an increasing function. Therefore, following the pioneering work of Dumortier and Roussarie [15], see also [30] by Krupa and Szmolyan, we can use the backward flow to extend the center manifold in the $\breve{v}=-1$-chart, see (5.21), into the scaling chart $\breve{\epsilon}=1$ (since $\rho_{3}=-p$ is decreasing) and continue this as a smooth invariant manifold there upon using the slow manifold, see (5.18), which is also attracting for the backward flow. Moreover, from here we can extend the manifold into the $\breve{v}=1$-chart by using the attracting center manifold, see (5.20). In this way, we obtain a global invariant manifold
@@ 5.1. Computing the difference
The proof is by contradiction: $\psi^{+}_{0}$ and $\psi^{-}_{0}$ satisfy the same differential equation on $x\in S^{+}\cap S^{-}$. This equation is clearly regular away from $x\neq 0$. Therefore if $\psi^{+}_{0}(x)=\psi^{-}_{0}(x)$ for some $x\in S^{+}\cap S^{-}$, $x\neq 0$, then by uniqueness of solutions, we have $\psi^{+}_{0}(x)=\psi^{-}_{0}(x)$ for all $x\in B_{\xi}$, i.e. in a full neighborhood of the origin. But then by Riemann’s theorem of removable singularities, we conclude that the unperturbed manifolds are analytic at the origin.
@@
with $\Delta s(-\delta)=\overline{\Theta_{u}}$ as the initial condition. Since the system is linear, we can integrate the equation directly and if $\Theta_{u}(0)\neq 0$ then we easily conclude that $\Delta s(0)$ is exponentially large with respect to $\epsilon\rightarrow 0$. But this contradicts Proposition 3.3 and the boundedness of $m_{2}^{\pm}(0,\epsilon)$. We therefore conclude that $\Theta_{s}(0)\neq 0$ and instead use $\Delta s(\delta)=\Theta_{s}$ as an initial condition for (5.25). We obtain

### 17 Apr 2026 v2

@@ 5. A dynamical systems approach to the difference
The system (5.14) can be viewed as a Fenichel normal form, see [28, Section 3]. The following result shows that the diagonalization (5.13) can be “globalized” beyond a compact subset $v_{2}\in I$ (by working in the directional charts $\breve{v}=\pm 1$ associated with the blowup (5.4)).
@@
Now, notice that with respect to the backward flow, the center manifolds in the $\breve{v}=\pm 1$-charts and the slow manifold in the $\breve{\epsilon}=1$-chart all become attracting and moreover $v$ (in particular $v_{2}$) becomes an increasing function. Therefore, following the pioneering work of Dumortier and Roussarie [18], see also [33] by Krupa and Szmolyan, we can use the backward flow to extend a fixed center manifold in the $\breve{v}=-1$-chart, see (5.21), into the scaling chart $\breve{\epsilon}=1$ (since $\rho_{3}=-p$ is decreasing) and continue this as a smooth invariant manifold there upon using the slow manifold, see (5.18), which is also attracting for the backward flow. Moreover, from here we can extend the resulting manifold into the $\breve{v}=1$-chart by using the attracting center manifold, see (5.20). In this way, we obtain a global invariant manifold
@@ 5.1. Computing the difference
The proof is by contradiction. First, we note that $\psi^{+}_{0}$ and $\psi^{-}_{0}$ satisfy the same differential equation on $x\in S^{+}\cap S^{-}$, recall Lemma 2.1. This equation is clearly regular away from $x\neq 0$. Therefore if $\psi^{+}_{0}(x)=\psi^{-}_{0}(x)$ for some $x\in S^{+}\cap S^{-}$, $x\neq 0$, then by uniqueness of solutions, we have $\psi^{+}_{0}(x)=\psi^{-}_{0}(x)$ for all $x\in B_{\xi}$, i.e. in a full neighborhood of the origin. But then by Riemann’s theorem of removable singularities, we conclude that the unperturbed manifolds are analytic at the origin.
@@
with $\Delta s(-\delta)=\overline{\Theta_{u}}$ as the initial condition. Since the system is linear, we can integrate the equation directly and if $\Theta_{u}(0)\neq 0$ then we easily conclude that $\Delta s(0)$ is exponentially large with respect to $\epsilon\rightarrow 0$ (see also Section 5.2 below). But this contradicts Proposition 3.3 and the boundedness of $m_{2}^{\pm}(0,\epsilon)$. We therefore conclude that $\Theta_{s}(0)\neq 0$ and instead use $\Delta s(\delta)=\Theta_{s}$ as an initial condition for (5.25). We obtain